# An LLM-Infused Tool for Workers’ Collective Intelligence

### v2

@@
Abstract.
Collective intelligence among gig workers yields considerable advantages, including improved information exchange, deeper social bonds, and stronger advocacy for better labor conditions. Especially as it enables workers to collaboratively pinpoint shared challenges and devise optimal strategies for addressing these issues. However, enabling collective intelligence remains challenging, as existing tools often overestimate gig workers’ available time and uniformity in analytical reasoning. To overcome this, we introduce GigSense, a tool that leverages large language models alongside theories of collective intelligence and sensemaking. GigSense enables gig workers to rapidly understand and address shared challenges effectively, irrespective of their diverse backgrounds. GigSense not only empowers gig workers but also opens up new possibilities for supporting workers more broadly, demonstrating the potential of large language model interfaces to enhance collective intelligence efforts in the evolving workplace.
Gig work, Crowdsourcing, Future of work, Generative AI, Human-AI Collaborations
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computingccs: Collaborative and social computing systems and tools
Figure 1. Overview of our process for designing and studying a generative AI tool for gig workers.
@@ 1. Introduction
Another hurdle is that the diverse and significant time commitments of gig workers may limit their ability to contribute to collective intelligence, underscoring the importance of developing inclusive tools that can accommodate workers’ varying schedules, reducing the barriers to collective intelligence and enabling easier participation in collaboratively tackling challenges (Packham, 2008; Pantea, 2013). To address these challenges, we introduce GigSense, a novel platform for fostering collective intelligence among gig workers by enabling collaborative problem-solving. By integrating Sense-Making Theory and leveraging the capabilities of large-language models (LLMs), GigSense features an LLM-enhanced interface that allows workers to deeply analyze their challenges and solutions from multiple perspectives (Peter and Card, 2005). Unlike existing platforms that offer a simple list-based view of issues (Salehi et al., 2015), GigSense enables a detailed exploration of problems, offering both a close-up and a broad overview to understand workplace dynamics better. Additionally, GigSense uses LLMs to facilitate collective brainstorming, helping workers develop solutions together. Fig 1 presents an overview of GigSense. As a significant advancement in fostering collective intelligence among gig workers, GigSense addresses the unique challenges of this workforce, which typically operates in isolation without a common platform for problem identification and solution generation (Calacci, 2022; Yao et al., 2021). GigSense bridges this gap by offering a communal space for gig workers to collectively understand and address their issue. Leveraging shared digital resources to build community (Sawyer et al., 2014; Barley, 1986), GigSense emerges as a vital tool for a fragmented workforce, potentially catalyzing the formation of supportive gig worker communities (Tajfel et al., 1979; Cohen et al., 2003; Van Zomeren et al., 2008; Tajfel and Turner, 2004). Beyond organization, it could reveal systemic issues, encouraging workers to see problems like client disputes and payment delays as collective challenges. This recognition can promote solidarity, crucial for collective intelligence, and empower workers to develop effective negotiation strategies, potentially improving their working conditions.
@@
In this paper, we highlight a system that integrates LLMs with an interactive interface to support gig workers in problem-solving and kickstart collective intelligence among workers.
@@ 2. GigSense
GigSense is an AI-enhanced platform designed to assist gig workers in the early phases of sensemaking and collective intelligence, which includes problem identification and solution proposal (Shaw et al., 2014). Our platform thus focuses on aiding workers in their sensemaking process (identifying collective issues and potential solutions). For this purpose, GigSense incorporates Pirolli and Card’s sensemaking theories and collective intelligence theories into various modules within its system (Shaw et al., 2014; Weick, 1995; Mulgan, 2018a). These modules are designed to automate the sensemaking process, which unfolds seamlessly within the entirety of GigSense’s interface, with workers playing a crucial role in steering this process.
@@
GigSense’s modules are based on previous designs (Flores-Saviaga et al., 2022; Li et al., 2018), but uniquely integrate human-AI collaboration to address specific needs and contexts of gig workers. GigSense recognizes the time constraints faced by gig workers when engaging in collective intelligence and hence provides interfaces that facilitate rapid sensemaking of problems and proposals of solutions. For this purpose, it utilizes Large Language Models (LLMs) to provide concise summaries of issues found on gig platforms, enabling workers to quickly understand prevalent problems. By leveraging the OpenAI GPT-4 API (OpenAI, 2024), it categorizes data and generates these summaries, aiding workers in gaining insights and brainstorming solutions. Our system also offers an interactive interface, allowing users to “zoom in” and “zoom out” and explore specific discussions related to a problem. While LLM supports the creation of problem summaries, GigSense’s interface design empowers workers to independently assess the information. It also provides a collaborative workspace for workers to propose solutions to identified problems. It uses LLMs to inspire workers with potential solutions but prioritizes human-generated ideas over AI suggestions to emphasize the value of human input. Notice that GigSense’s design, focusing on quick sensemaking, contrasts with traditional extended analysis but offers notable advantages. It enables time-constrained individuals like gig workers with multiple jobs, parents, and students to participate, enhancing diversity and inclusivity. This approach makes sensemaking more accessible, as it requires less long-term commitment (Etz et al., 2018). A faster process encourages broader participation and engagement, as people are more likely to contribute when their time investment is minimal yet impactful (Sandberg and Tsoukas, 2015). This approach can be especially beneficial for future engagements around collective intelligence, such as implementing solutions, where sustained participant involvement is crucial (Etz et al., 2018). Additionally, rapid sensemaking can foster concise and focused information processing, reducing data overload and maintaining participant interest, which can wane in longer processes (Mulgan, 2018b; Sandberg and Tsoukas, 2015). Next, we describe each GigSense module in detail:
@@
Gig workers supply Gigsense with a roster of subreddits from which they intend to pinpoint potential problems and datasets containing assessments for gig work platforms (This is the “Step: Search and Filter” in Pirolli et al.’s sensemaking loop (Peter and Card, 2005)). Next, Gigsense connects to the Reddit API to read and extract all the posts from the subreddits that gig workers initially provide. GigSense additionally uses a web scrapper to extract data from reviews left on Apple’s and Google’s app stores by gig workers. Note that our data gathering module only collects reviews that have between one and three-star ratings. The module considers that these review data would represent complaints and problems that gig workers are experiencing. Gigsense also lets workers manually enter issues into the system if they choose to do so (“Step: Read and Extract” in the sensemaking loop). Using the real-world gig workers’ complaint datasets (actual gig workers’ subreddits and complaints) in our system design aims to bring inclusiveness about gig worker concerns and complaints. The show button allows workers to see all the data sources used in GigSense.
@@
Acknowledging the potential enormity of the data gathered through the Data Gathering Module and its potential complexity for human interpretation, this module centers its efforts on summarizing the data (“Step: Schematize” in the sensemaking loop). For this purpose, Gigsense uses LLMs to facilitate the exploration and sensemaking of the extensive datasets of complaints, enabling workers to categorize and summarize the large datasets into problem categories. This module also has two buttons to allow workers to navigate to the Data Visualization module, which offers a “zoom-in” view for meticulous examination of the data or a “zoom-out” view for a panoramic grasp of their issues.
@@
This module enables workers to analyze problems at various levels, aiding their sensemaking process. It utilizes semantic zooming, a method for visualizing data across different abstraction levels (Bederson and Hollan, 1994; Suh et al., 2023a, b). Workers have access to two views showcasing varying detail levels for problem analysis. The zoomed-out view of this module presents an interactive chart of the problems (“Step: Schematize” in the sensemaking loop) where workers can grasp a high-level understanding of the specific types of problems (e.g. Payment, Platform Policy, Scam, Customer Support, Usability, etc) faced by them. Gigsense utilizes LLMs to streamline the sensemaking process by categorizing dataset posts into distinct categories. These categorized posts are later used by the system to create the different graphs and the corresponding category descriptions. The use of LLMs speeds up data organization, allowing workers to engage in sensemaking more efficiently. The interactive chart also serves as a powerful tool for raising awareness among gig workers about their shared challenges. For instance, workers can engage with these graphs by hovering over them, enabling access to the exact number posts (complaints) within each category. This functionality contributes to the worker’s sensemaking process by facilitating the easy identification of prevalent issues.
The volume of information available to workers shifts according to the zoom level. To let the worker delve deeper into individual problems, Gigsense offers a zoom-in perspective. Choosing the zoom-in view shifts the display from graphs to the full text of posts relevant to the problem being analyzed. This view allows workers to view the nuanced aspects of issues that pertain to the topic analyzed and upvote on individual problems. By visually displaying the prevalence of issues (number of upvotes), it creates a sense of solidarity (Savage and Jarrahi, 2020). Workers can see that they are not alone in their struggles, fostering a collective identity and shared purpose (Lee et al., 2015). It provides a static depiction of the number of workers grappling with problems across different gig work platforms, acting as a catalyst in building communities of gig workers facing similar challenges to encourage collective intelligence. The zoomable interface facilitates sensemaking by allowing users to toggle between high-level context and detailed analysis. Note that as (Mulgan, 2018b) discusses, collective intelligence requires focused attention on what matters. The system is thus designed to allow workers to analyze problems from detailed and broader perspectives.
@@
This module further facilitates the sensemaking process and focuses on helping gig workers to develop concrete solutions to address the problem analyzed (“Step: Build Case’ ” in the sensemaking loop). It incorporates sub-modules such as the: “Sensemaking Chat”, “Shared Document” and “Collaborative Solution Space”. The Sensemaking Chat submodule allows workers to engage in conversations to discuss and investigate the problems they encounter in their work. They can communicate through asynchronous text messages to accommodate different schedules. The “Shared Document” enables workers to understand the existing problems and collaboratively create action plans (solutions) to address them. Gigsense also includes a functionality that allows users to make annotations in the shared document. This way, workers can collectively review and approve their proposed solutions. Finally, the “Collaborative Solution Space”, just like the sensemaking process, features a space where workers can showcase the final solution they mutually agreed upon (Weick, 1995; Dervin, 1983) (“Step: Tell Story” in the sense-making loop). This module supports collaborative work that initiates early stages of collective intelligence(Shaw et al., 2014).

### v3

# GigSense: An LLM-Infused Tool for Workers’ Collective Intelligence

@@
Abstract
Collective intelligence among gig workers yields considerable advantages, including improved information exchange, deeper social bonds, and stronger advocacy for better labor conditions. Especially as it enables workers to collaboratively pinpoint shared challenges and devise optimal strategies for addressing these issues. However, enabling collective intelligence remains challenging, as existing tools often overestimate gig workers’ available time and uniformity in analytical reasoning. To overcome this, we introduce GigSense, a tool that leverages large language models alongside theories of collective intelligence and sensemaking. GigSense enables gig workers to rapidly understand and address shared challenges effectively, irrespective of their diverse backgrounds. Our user study showed that GigSense users outperformed those using a control interface in problem identification and generated solutions more quickly and of higher quality, with better usability experiences reported. GigSense not only empowers gig workers but also opens up new possibilities for supporting workers more broadly, demonstrating the potential of large language model interfaces to enhance collective intelligence efforts in the evolving workplace.
@@ 1. Introduction
Another hurdle is that the diverse and significant time commitments of gig workers may limit their ability to contribute to collective intelligence, underscoring the importance of developing inclusive tools that can accommodate workers’ varying schedules, reducing the barriers to collective intelligence and enabling easier participation in collaboratively tackling challenges (Packham, 2008; Pantea, 2013). To address these challenges, we introduce GigSense, a novel platform for fostering collective intelligence among gig workers by enabling collaborative problem-solving. By integrating Sense-Making Theory and leveraging the capabilities of large-language models (LLMs), GigSense features an LLM-enhanced interface that allows workers to deeply analyze their challenges and solutions from multiple perspectives (Peter and Card, 2005). Unlike existing platforms that offer a simple list-based view of issues (Salehi et al., 2015), GigSense enables a detailed exploration of problems, offering both a close-up and a broad overview to understand workplace dynamics better. Additionally, GigSense uses LLMs to facilitate collective brainstorming, helping workers develop solutions together. Our user study indicates that GigSense enhances the sensemaking process for gig workers, accelerating their ability to identify a broader range of problems and generate more feasible, diverse solutions. As a significant advancement in fostering collective intelligence among gig workers, GigSense addresses the unique challenges of this workforce, which typically operates in isolation without a common platform for problem identification and solution generation (Calacci, 2022; Yao et al., 2021). GigSense bridges this gap by offering a communal space for gig workers to collectively understand and address their issue. Leveraging shared digital resources to build community (Sawyer et al., 2014; Barley, 1986), GigSense emerges as a vital tool for a fragmented workforce, potentially catalyzing the formation of supportive gig worker communities (Tajfel et al., 1979; Cohen et al., 2003; Van Zomeren et al., 2008; Tajfel and Turner, 2004). Beyond organization, it could reveal systemic issues, encouraging workers to see problems like client disputes and payment delays as collective challenges. This recognition can promote solidarity, crucial for collective intelligence, and empower workers to develop effective negotiation strategies, potentially improving their working conditions.
@@
In this paper, we highlight two main contributions: (1) a system that integrates LLMs with an interactive interface to support gig workers in problem-solving and kickstart collective intelligence among workers, and (2) a study showing GigSense not only speeds up problem-solving but also improves the quality and variety of worker-generated solutions. We discuss the influence of LLMs on sensemaking and collective intelligence and its implications for workers, proposing directions for future research.
2. Related Work
Gig Work and Platform
“Gig” or platform-based work is a significant trend in the labor market, driven by the demand for flexibility from both employers and workers (Watson et al., 2021), and facilitated by digital technology (Mandl et al., 2015; De Stefano, 2016; Johnston and Land-Kazlauskas, 2018). While offering economic benefits to disadvantaged groups, gig work also presents challenges like unstable schedules, income variability, and uncertain long-term job security (Peck and Theodore, 2012; De Stefano, 2016). Workers have engaged in collective actions (e.g., negotiations, strikes, unionization) to improve conditions, yet face obstacles due to platform constraints, geographic dispersion, and a lack of community and common interest, making organizing difficult (Gray and Suri, 2019; Johnston and Land-Kazlauskas, 2018; Sutherland and Jarrahi, 2017; Yao et al., 2021; Webster, 2016). Despite these challenges, some success has been noted in collective efforts (Dubal, 2017), but tools to support organizing are limited (Johnston and Land-Kazlauskas, 2018). This paper proposes a tool to aid gig workers in understanding their challenges and initiating collective intelligence.
Sensemaking and Collective Intelligence
Comprehending the challenges faced by gig workers can be perceived as an act of sensemaking, involving the collection and analysis of diverse and unstructured data to reach a conclusion. Pirolli and Card (Peter and Card, 2005) define sensemaking as a series of iterative steps. For instance, it starts with the initial gathering of relevant data (“Step: Search and Filter”), akin to brainstorming gig workers’ problems. Subsequently, it involves extracting valuable information (“Step: Read and Extract”), akin to selecting the most pertinent issues. Further, it encompasses summarizing and schematizing the information (“Step: Schematize”), akin to the manual procedure of condensing and structuring of the brainstormed ideas. Then, it involves generating hypotheses from various perspectives (“Step: Build Case”), resembling the development of viable solutions. Lastly, it culminates in decision-making to determine the best solution (“Step: Tell Story”). Smith (Smith, 1994) defines collective intelligence as a group of individuals working together on tasks, where the collective group itself demonstrates coherence, intelligence, and constant improvement, enabling more effective mobilization of skills than any single individual working independently. Significant research explores collaborative sensemaking tools for domains like literature review (Zhang et al., 2008), web search (Paul and Morris, 2009), organizing academic literature (Reimer et al., 2011), solving mysteries (Li et al., 2018), and tackling disinformation (Flores-Saviaga et al., 2022). We introduce GigSense to aid gig workers’ collaborative sensemaking, automating parts of the process pipeline. We also leverage Large Language Models to enhance collective intelligence.
LLMs for Idea Generation.
In recent years, significant advancements in large language models (LLMs) have positioned them as a promising tool for facilitating a diverse array of writing tasks, such as story generation (Mirowski et al., 2023; Chung et al., 2022; Yuan et al., 2022; Alabdulkarim et al., 2021), academic writing (Gero et al., 2022), question-answering (Brown et al., 2020), idea generation (Girotra et al., 2023) and information synthesis (Liu et al., 2023; Palani et al., 2023). However, despite their impressive utility, LLMs have faced criticism for generating text that, while appearing logically and grammatically coherent, may contain factual inaccuracies or lack meaningful coherence (referred to as hallucinations) (Ji et al., 2023). Nevertheless, researchers suggest that this weakness can be reframed as a strength (Girotra et al., 2023), and have started to use LLMs for idea suggestions(Yuan et al., 2022). In creative writing and problem-solving, possessing a range of idea quality and quantity holds more value than maintaining unwavering consistency (Girotra et al., 2010). To achieve a wide range of ideas with varying levels of quality, most ideation research advises generating numerous ideas first and delaying their evaluation (Girotra et al., 2010). LLMs are designed to do exactly this— quickly generate many somewhat plausible solutions (Girotra et al., 2023). In this research, we align with the principle that an unrestrained influx of ideas can often pave the way for innovative solutions. Therefore, we harness the capabilities of LLMs by integrating them into Gigsense to assist gig workers in rapidly producing multiple reasonably viable ideas for their most pressing challenges.
Interfaces for Visualizing Collective Problems and Solutions
. Other research that has inspired our work includes interfaces and systems designed to assist individuals in visualizing both problems and solutions. For example, MacNeil et al. (MacNeil et al., 2021), presented a design gallery for visualizing problems and the stakeholders involved in those problems. Similarly, Huang et al. (Huang and MacNeil, 2023) introduced a novel system to aid designers in comprehending various issues within a space and exploring diverse solutions. Their research focused on mitigating design fixation by encouraging a broader consideration of context and essential relationships during the design process. Siangliulue et al.’s system (Siangliulue et al., 2016) innovatively merged crowdsourcing with machine learning to create a semantic solution space model, whihelped to foster user creativity and diversity of ideas. Our research is inspired by the principles of these systems, concentrating on assisting gig workers, who may lack expertise in design and problem-solving, in exploring and comprehending various aspects of problems and having support to provide solutions.
@@ 3. GigSense
Diagram of GigSense’s problem-solving interface showing four modules: A) Data Gathering Module; B) Collaborative Solution Module where workers can collaborate on solutions to specific problems; C) Problem Summary Module where workers get summaries of specific problems to aid sensemaking; D) AI-Enhanced Solution Module where workers get solution suggestions from large language models.
Figure 2. Overview of GigSense’s problem-solving interface with the: A) “Data Gathering Module”; B) “Collaborative Solution Module” where workers can collaborate to create solutions to specific problems; C) “Problem Summary Module” where workers obtain summaries of specific problems to help their sensemaking; and D) “AI-Enhanced Solution Module” where workers obtain solution suggestions from LLMs.
3.1. Description of GigSense
@@
GigSense is an AI-enhanced platform designed to assist gig workers in the early phases of sensemaking and collective intelligence, which includes problem identification and solution proposal (Shaw et al., 2014). Our platform thus focuses on aiding workers in their sensemaking process (identifying collective issues and potential solutions). For this purpose, GigSense incorporates Pirolli and Card’s sensemaking theories and collective intelligence theories into various modules within its system (Shaw et al., 2014; Weick, 1995; Mulgan, 2018a). Fig 2 presents GigSense’s interface and its modules. These modules are designed to automate the sensemaking process, which unfolds seamlessly within the entirety of GigSense’s interface, with workers playing a crucial role in steering this process.
@@
GigSense’s modules are based on previous designs (Flores-Saviaga et al., 2022; Li et al., 2018), but uniquely integrate human-AI collaboration to address specific needs and contexts of gig workers (see Fig 2.C-D.) GigSense recognizes the time constraints faced by gig workers when engaging in collective intelligence and hence provides interfaces that facilitate rapid sensemaking of problems and proposals of solutions. For this purpose, it utilizes Large Language Models (LLMs) to provide concise summaries of issues found on gig platforms, enabling workers to quickly understand prevalent problems. By leveraging the OpenAI GPT-4 API (OpenAI, 2024), it categorizes data and generates these summaries, aiding workers in gaining insights and brainstorming solutions. Our system also offers an interactive interface, allowing users to “zoom in” and “zoom out” and explore specific discussions related to a problem. While LLM supports the creation of problem summaries, GigSense’s interface design empowers workers to independently assess the information. It also provides a collaborative workspace for workers to propose solutions to identified problems. It uses LLMs to inspire workers with potential solutions but prioritizes human-generated ideas over AI suggestions to emphasize the value of human input. Notice that GigSense’s design, focusing on quick sensemaking, contrasts with traditional extended analysis but offers notable advantages. It enables time-constrained individuals like gig workers with multiple jobs, parents, and students to participate, enhancing diversity and inclusivity. This approach makes sensemaking more accessible, as it requires less long-term commitment (Etz et al., 2018). A faster process encourages broader participation and engagement, as people are more likely to contribute when their time investment is minimal yet impactful (Sandberg and Tsoukas, 2015). This approach can be especially beneficial for future engagements around collective intelligence, such as implementing solutions, where sustained participant involvement is crucial (Etz et al., 2018). Additionally, rapid sensemaking can foster concise and focused information processing, reducing data overload and maintaining participant interest, which can wane in longer processes (Mulgan, 2018b; Sandberg and Tsoukas, 2015). Next, we describe each GigSense module in detail:
Diagram showing GigSense’s zoom-out and zoom-in views accessible via interface buttons. Zoom-out shows an overall view of the problems. Zoom-in shows a concise summary of specific problems being analyzed and relevant posts.
Figure 3. Gigsense’s zoom-out and zoom-in view. These views can be accessed by gig workers using the corresponding buttons in the main interface. Gig-sense Zoomed-out view presents an overall view of the problems. Gig-sense Zoom-in view delivers a concise summary of the problem being analyzed, while also enabling users to explore relevant posts associated with that specific challenge.
@@
Gig workers supply Gigsense with a roster of subreddits from which they intend to pinpoint potential problems and datasets containing assessments for gig work platforms (This is the “Step: Search and Filter” in Pirolli et al.’s sensemaking loop (Peter and Card, 2005)). Next, Gigsense connects to the Reddit API to read and extract all the posts from the subreddits that gig workers initially provide. GigSense additionally uses a web scrapper to extract data from reviews left on Apple’s and Google’s app stores by gig workers. Note that our data gathering module only collects reviews that have between one and three-star ratings. The module considers that these review data would represent complaints and problems that gig workers are experiencing. Gigsense also lets workers manually enter issues into the system if they choose to do so (“Step: Read and Extract” in the sensemaking loop). Using the real-world gig workers’ complaint datasets (actual gig workers’ subreddits and complaints) in our system design aims to bring inclusiveness about gig worker concerns and complaints. The show button allows workers to see a list of all the data sources used in GigSense.
@@
Acknowledging the potential enormity of the data gathered through the Data Gathering Module and its potential complexity for human interpretation, this module centers its efforts on summarizing the data (“Step: Schematize” in the sensemaking loop). For this purpose, Gigsense uses LLMs to facilitate the exploration and sensemaking of the extensive datasets of complaints, enabling workers to categorize and summarize the large datasets into problem categories. This module also has two buttons to allow workers to navigate to the Data Visualization module, which offers a “zoom-in” view for meticulous examination of the data or a “zoom-out” view for a panoramic grasp of their issues (See Fig 2.C).
@@
This module enables workers to analyze problems at various levels, aiding their sensemaking process. It utilizes semantic zooming, a method for visualizing data across different abstraction levels (Bederson and Hollan, 1994; Suh et al., 2023a, b). Workers have access to two views (See Fig.3) showcasing varying detail levels for problem analysis. The zoomed-out view of this module presents an interactive chart of the problems (“Step: Schematize” in the sensemaking loop) where workers can grasp a high-level understanding of the specific types of problems (e.g. Payment, Platform Policy, Scam, Customer Support, Usability, etc) faced by them. Gigsense utilizes LLMs to streamline the sensemaking process by categorizing dataset posts into distinct categories. These categorized posts are later used by the system to create the different graphs and the corresponding category descriptions. The use of LLMs speeds up data organization, allowing workers to engage in sensemaking more efficiently. The interactive chart also serves as a powerful tool for raising awareness among gig workers about their shared challenges. For instance, workers can engage with these graphs by hovering over them, enabling access to the exact number posts (complaints) within each category. This functionality contributes to the worker’s sensemaking process by facilitating the easy identification of prevalent issues commonly cited in complaints.
The volume of information available to workers shifts according to the zoom level. To let the worker delve deeper into individual problems, Gigsense offers a zoom-in perspective. Choosing the zoom-in view shifts the display from graphs to the full text of posts relevant to the problem being analyzed. This view allows workers to view the nuanced aspects of issues that pertain to the topic analyzed and upvote on individual problems. By visually displaying the prevalence of issues (number of upvotes), it creates a sense of solidarity (Savage and Jarrahi, 2020). Workers can see that they are not alone in their struggles, fostering a collective identity and shared purpose (Lee et al., 2015). It provides a static depiction of the number of workers grappling with problems across different gig work platforms, acting as a catalyst in building communities of gig workers facing similar challenges to encourage collective intelligence. The zoomable interface facilitates sensemaking by allowing users to toggle between high-level context and detailed analysis. Note that as (Mulgan, 2018b) discusses, collective intelligence requires focused attention on what matters. The system is thus designed to aid workers in focusing on their problems, allowing them to analyze them from both detailed and broader perspectives.
@@
This module further facilitates the sensemaking process and focuses on helping gig workers to develop concrete solutions to address the problem analyzed (“Step: Build Case’ ” in the sensemaking loop). It incorporates sub-modules such as the: “Sensemaking Chat”, “Shared Document” and “Collaborative Solution Space”. The Sensemaking Chat submodule allows workers to engage in conversations to discuss and investigate the problems they encounter in their work. They can communicate through asynchronous text messages to accommodate different schedules. The “Shared Document” enables workers to understand the existing problems and collaboratively create action plans (solutions) to address them. Gigsense also includes a functionality that allows users to make annotations in the shared document. This way, workers can collectively review and approve their proposed solutions. Finally, the “Collaborative Solution Space”, just like the sensemaking process, features a space where workers can showcase the final solution they mutually agreed upon (Weick, 1995; Dervin, 1983) (“Step: Tell Story” in the sense-making loop).Note that this module supports collaborative work that initiates early stages of collective intelligence(Shaw et al., 2014).